-5/8 divided by 3/4
Found 2 solutions by CassieLynn, rfer:
Answer by CassieLynn(3)

-5/8=-0.62
3/4=0.75
-0.62/0.75=-0.83

Answer by rfer(2688)

(-5/8)/(3/4)
invert and multiply
-5/8*4/3=-20/24=-5/6